# Generalized Heisenberg Dynamics Revisited

### v1

@@ 3.2.1 Fermionic harmonic oscillator
We study a harmonic oscillator relating to fermionic variables as an example of a solution describing eqs. (84) and (85). They are represented by the $3\times 3\times 3$ cubic matrices whose components are given by
@@ 4.1 Generalized matrix mechanics
where sgn(P) is $+1$ and $-1$ for even and odd permutations among the set of subscripts $(a_{1},a_{2},\cdots,a_{n-1})$, respectively, based on $+1$ for $(a_{1},a_{2},\cdots,a_{n-1})=(1,2,\cdots,n-1)$, and each $a_{k}$ $(k=1,2,\cdots,n-1)$ runs from $1$ to $n-1$. In the following, we construct a mechanics of quantities with $n$ indices, using the correspondence principle, or tackle a puzzle like a cryptarithm shown in Figure 3.
@@
where $\beta$ is a constant, and $(\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n}}$ is regarded as a counterpart of $\nu\varDelta l_{1}\varDelta l_{2}\cdots\varDelta l_{n-1}$. From eq. (108), we see that $(\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots l_{n}}$ is anti-symmetric under the exchange of $n-1$ indices $(l_{1},l_{2},\cdots,l_{n-1})$.
Using eqs. (106), (107) and (108), we derive the relation666 The details of derivation of this relation and most other equations are presented in Appendix A.:
@@
Let us require that $(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ should have a cyclic property among indices777 The frequency $(\nu_{n}^{(0)})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ given by eq. (108) also have a cyclic property in a hidden way, as shown in Appendix A, and hence we find that it is totally anti-symmetric under the exchange of $(l_{1},l_{2},\cdots,l_{n})$. such as
@@
If we require that the product $(A_{1}A_{2}\cdots A_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)$ should take the same form as eq. (115), i.e., $(A_{1}A_{2}\cdots A_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)=(A_{1}A_{2}\cdots A_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(0)e^{2\pi i(\nu_{n})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}t}$, $(A_{a})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)$ become components of extended matrices whose product is defined by888 This definition of the $n$-fold product is the same as that in Ref. [22].
@@
where $c_{l_{u}~{}\!\!l_{i}}$ are constants, $i$ and $j(\neq i)$ are some numbers from $1$ to $n$, $u$ runs from 1 to $n$ except for $i$ and $j$, and the $n-1$ indices $(l_{1},\cdots,l_{i},\cdots l_{j-1},l_{j+1},\cdots,l_{n})$ are all different. We note that other components are zero. We refer to this type of generalized matrix as a normal form, a normal generalized matrix or an $n$-th power normal matrix.
@@
where $(h_{a_{k}})_{l_{u}~{}\!\!l_{i}}$ are constants, $i$ and $j(\neq i)$ are some numbers from $1$ to $n$, $u$ runs from 1 to $n$ except for $i$ and $j$, and the $n-1$ indices $(l_{1},\cdots,l_{i},\cdots l_{j-1},l_{j+1},\cdots,l_{n})$ are all different.
@@
×$~{}~{}~{}~{}\times\prod_{r=1}^{k-1}(E_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(E_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}A_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)$
@@
×$~{}~{}~{}~{}\times\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}.$ (123)
@@
×$~{}~{}~{}~{}~{}~{}~{}\left.\times(H_{a_{3}})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n-3}~{}\!\!l_{n+1}~{}\!\!l_{n-1}~{}\!\!l_{n}}\cdots(H_{a_{n-1}})_{l_{1}~{}\!\!l_{n+1}~{}\!\!l_{3}~{}\!\!\cdots~{}\!\!l_{n}}A(t)_{l_{n+1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}\right\}.$
@@
In the same way as the dynamics of cubic matrices, it is reconfirmed that the components with identical indices on $A_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}(t)$ are not independent of time, using eq. (125) and the relation $[B,C_{1},\cdots,C_{n-2},C_{n-1}]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}=0$ that holds concerning a generalized matrix $B$ whose components with all different indices vanish, i.e., $B_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}=0$ ($l_{i}\neq l_{j}$ for all $i$ and $j$) and arbitrary normal cubic matrices $C_{a}$ $(a=1,2,\cdots,n-1)$. The validity of $[B,C_{1},\cdots,C_{n-2},C_{n-1}]_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}=0$ is understood from the following reason. For $B$ and $C_{a}$, we obtain the relation:
@@
×$~{}~{}~{}~{}~{}\times(C_{n-2})_{l_{1}~{}\!\!l_{n+1}~{}\!\!l_{3}~{}\!\!\cdots~{}\!\!l_{n}}(C_{n-1})_{l_{n+1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}=0$ (126)
from the fact that there exists no instance in which each of indices $(l_{1},l_{2},\cdots,l_{n-2},l_{n+1},l_{n})$, $\cdots$, $(l_{1},l_{n+1},l_{3},\cdots,l_{n})$ and $(l_{n+1},l_{2},\cdots,l_{n})$ contains exactly two identical indices, other indices are all different, $(l_{1},l_{2},\cdots,l_{n-1},l_{n+1})$ contains two or more than two identical indices, and furthermore, the identical index is shared between them.
@@
One is a symmetry in the equation. Eq. (125) is invariant under the transformation of Hamiltonians $(H_{a_{k}})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}\to(H^{\prime}_{a_{k}})_{l_{1}~{}\!\!l_{2}~{}\!\!\cdots~{}\!\!l_{n}}$ whose values given by eq. (118) are transformed as
@@
×$~{}~{}~{}~{}\times\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-2}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}$ (133)
@@
The other is a solution of eq. (125), in the case that eq. (125) is regarded as a master equation. There is another solution such that
@@
×$~{}~{}~{}~{}\times\prod_{r=1}^{k-1}(\tilde{h}_{a_{r}})_{l_{r}~{}\!\!l_{k}}\prod_{r^{\prime}=k}^{n-2}(\tilde{h}_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~{}\!\!l_{k}}$ (137)
@@ 4.2 Discretization of Jacobian
where sgn(P) is $+1$ and $-1$ for even and odd permutations among the set of subscripts $(a_{1},a_{2},\cdots,a_{n})$, respectively, based on $+1$ for $(a_{1},a_{2},\cdots,a_{n})=(1,2,\cdots,n)$, and each $a_{k}$ $(k=1,2,\cdots,n)$ runs from $1$ to $n$. Hence, we conjecture that the generalized Heisenberg’s equation of motion (125) can be a counterpart of the Nambu equation and the multiple commutator defined by
@@ 5 Conclusions and discussions
Inspired by the fact that Heisenberg’s matrix mechanics was derived from Hamiltonian mechanics under the guidance of the correspondence principle, we have investigated a mechanics involving discrete variables, starting from Nambu mechanics. As a result, we reconstructed a new framework of mechanics for dynamical variables expressed in terms of generalized matrices, which can be regarded as an extension of matrix mechanics. The fundamental equation of this mechanics corresponds to a generalized version of Heisenberg’s equation of motion, formulated using an $n$-fold commutator. This equation can also be interpreted as the counterpart to the Nambu equation associated with canonical $n$-plets. Based on this interpretation, we reconfirmed that the multiple commutator involving generalized matrices can serve as a quantum or discrete version of the Nambu bracket, namely the Jacobian.
@@
In quantum mechanics, the matrix elements in matrix mechanics are interpreted as probability amplitudes representing transitions between quantum states. At present, it is unclear whether the generalized matrix elements with $n$ indices admit a similar interpretation, or whether the extended version of matrix mechanics can be applied to real physical systems. This suggests the need for further investigation into the understanding and application of generalized matrices, beyond the framework of a mere extension of Heisenberg’s matrix mechanics or a counterpart to Nambu mechanics.
@@
Finally, generalized matrices may prove useful in formulating the quantum theory of extended objects. The action integral in string theory includes the area of the worldsheet, and this area is expressed using a two-dimensional Jacobian [26, 27]. Similarly, $p$-dimensional extended objects are described using $(p+1)$-dimensional Jacobians. By discretizing the Jacobian, it may become possible to microscopically formulate various branes including membrane [28], which could, in turn, be applicable to the formulation of M-theory999 Using Lie 3-algebra [29], an effective theory for multiple M2 branes called the BLG model has been constructed [30, 31, 32, 33]. See Ref. [34] for extensive reviews of $n$-nary algebra and its applications.. Recently, there have been attempts relating to many-index objects, for instance, to regularize the Nambu brackets using $2n$-index objects [35] and to construct gauge theories using a new type of non-associative algebra that includes both ordinary matrices and cubic matrices [36]. These developments suggest that there may still be room for various extensions and applications.

### v2

@@ 3.2.1 Fermionic harmonic oscillator
We study a harmonic oscillator relating to fermionic variables [11] as an example of a solution describing eqs. (93) and (94). They are represented by the $3\times 3\times 3$ cubic matrices whose components are given by
@@ 4.1 Generalized matrix mechanics
where sgn(P) is +1+1 and 1-1 for even and odd permutations among the set of subscripts $(a_{1},a_{2},\cdots,a_{n-1})$, respectively, based on +1+1 for $(a_{1},a_{2},\cdots,a_{n-1})=(1,2,\cdots,n-1)$, and each $a_{k}$ $(k=1,2,\cdots,n-1)$ runs from $1$ to n1n-1. The above eq. (114) can be derived from Nambu dynamics describing a specific periodic motion by the canonical $n$-plets ($\theta$, $J_{1}$, $\cdots$, $J_{n-1}$) including a cyclic coordinate $\theta$ and the Hamiltonians ($H_{1}$, $H_{2}$, $\cdots$, $H_{n-1}$). Then, as the Hamiltonians do not contain $\theta$, the equations of motion are given by
$\frac{d\theta}{dt}=\frac{\partial(H_{1},H_{2},\cdots,H_{n-1})}{\partial(J_{1},J_{2},\cdots,J_{n-1})},$~~ (115)
$\frac{dJ_{1}}{dt}=\frac{\partial(H_{1},\cdots,H_{n-2},H_{n-1})}{\partial(J_{2},\cdots,J_{n-1},\theta)}=0,~\cdots,~\frac{dJ_{n-1}}{dt}=\frac{\partial(H_{1},H_{2},\cdots,H_{n-1})}{\partial(\theta,J_{1},\cdots,J_{n-2})}=0.$ (116)
From eqs. (115) and (116), we find that $J_{k}$s become constants and then $\theta=\theta(t)$ is given by $\theta(t)=\nu t+\theta_{0}$, where $\theta_{0}$ is a constant and $\nu$ is also a constant given by eq. (114) relating to $E_{k}$s (the values of $H_{k}$s), respectively.
In the following, we construct a mechanics of quantities with $n$ indices, using the correspondence principle, or tackle a puzzle like a cryptarithm shown in Figure 3.
@@
where $\beta$ is a constant, and $(\nu_{n}^{(0)})_{l_{1}~\!\!l_{2}~\!\!\cdots l_{n}}$ is regarded as a counterpart of $\nu\varDelta l_{1}\varDelta l_{2}\cdots\varDelta l_{n-1}$. From eq. (119), we see that $(\nu_{n}^{(0)})_{l_{1}~\!\!l_{2}~\!\!\cdots l_{n}}$ is anti-symmetric under the exchange of n1n-1 indices $(l_{1},l_{2},\cdots,l_{n-1})$.
Using eqs. (117), (118) and (119), we derive the relation [^6] The details of derivation of this relation and most other equations are presented in Appendix A.:
@@
Let us require that $(\nu_{n})_{l_{1}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}}$ should have a cyclic property among indices [^7] The frequency $(\nu_{n}^{(0)})_{l_{1}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}}$ given by eq. (119) also have a cyclic property in a hidden way, as shown in Appendix A, and hence we find that it is totally anti-symmetric under the exchange of $(l_{1},l_{2},\cdots,l_{n})$. such as
@@
If we require that the product $(A_{1}A_{2}\cdots A_{n})_{l_{1}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}}(t)$ should take the same form as eq. (126), i.e., $(A_{1}A_{2}\cdots A_{n})_{l_{1}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}}(t)=(A_{1}A_{2}\cdots A_{n})_{l_{1}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}}(0)e^{2\pi i(\nu_{n})_{l_{1}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}}t}$, $(A_{a})_{l_{1}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}}(t)$ become components of extended matrices whose product is defined by [^8] This definition of the $n$-fold product is the same as that in Ref. [22].
@@
where $c_{l_{u}~\!\!l_{i}}$ are constants, $i$ and $j(\neq i)$ are some numbers from $1$ to $n$, $u$ runs from 1 to $n$ except for $i$ and $j$, and the n1n-1 indices $(l_{1},\cdots,l_{i},\cdots l_{j-1},l_{j+1},\cdots,l_{n})$ are all different. We note that other components are zero. We refer to this type of generalized matrix as a normal form, a normal generalized matrix or an $n$-th power normal matrix.
@@
where $(h_{a_{k}})_{l_{u}~\!\!l_{i}}$ are constants, $i$ and $j(\neq i)$ are some numbers from $1$ to $n$, $u$ runs from 1 to $n$ except for $i$ and $j$, and the n1n-1 indices $(l_{1},\cdots,l_{i},\cdots l_{j-1},l_{j+1},\cdots,l_{n})$ are all different.
@@
$~~~~\times\prod_{r=1}^{k-1}(E_{a_{r}})_{l_{r}~\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(E_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~\!\!l_{k}}A_{l_{1}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}}(t)$
@@
$~~~~\times\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~\!\!l_{k}}\prod_{r^{\prime}=k}^{n-1}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~\!\!l_{k}}.$ (134)
@@
$~~~~~~~\left.\times(H_{a_{3}})_{l_{1}~\!\!l_{2}~\!\!\cdots~\!\!l_{n-3}~\!\!l_{n+1}~\!\!l_{n-1}~\!\!l_{n}}\cdots(H_{a_{n-1}})_{l_{1}~\!\!l_{n+1}~\!\!l_{3}~\!\!\cdots~\!\!l_{n}}A(t)_{l_{n+1}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}}\right\}.$
@@
In the same way as the dynamics of cubic matrices, it is reconfirmed that the components with identical indices on $A_{l_{1}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}}(t)$ are independent of time, using eq. (136) and the relation $B,C_{1},\cdots,C_{n-2},C_{n-1}]_{l_{1}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}}=0$ that holds concerning a generalized matrix $B$ whose components with all different indices vanish, i.e., $B_{l_{1}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}}=0$ ($l_{i}\neq l_{j}$ for all $i$ and $j$) and arbitrary normal cubic matrices $C_{a}$ $(a=1,2,\cdots,n-1)$. The validity of $B,C_{1},\cdots,C_{n-2},C_{n-1}]_{l_{1}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}}=0$ is understood from the following reason. For $B$ and $C_{a}$, we obtain the relation:
@@
$~~~~~\times(C_{n-2})_{l_{1}~\!\!l_{n+1}~\!\!l_{3}~\!\!\cdots~\!\!l_{n}}(C_{n-1})_{l_{n+1}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}}=0$ (137)
from the fact that there exists no instance in which each of indices $(l_{1},l_{2},\cdots,l_{n-2},l_{n+1},l_{n})$, $\cdots$, $(l_{1},l_{n+1},l_{3},\cdots,l_{n})$ and $(l_{n+1},l_{2},\cdots,l_{n})$ contains exactly two identical indices, other indices are all different, $(l_{1},l_{2},\cdots,l_{n-1},l_{n+1})$ contains two or more than two identical indices, and furthermore, the identical index is shared between them. The equation (136) coincides with the generalized Heisenberg equation obtained in Ref. [14] by employing the generalized Ritz rule and the generalized Bohr frequency condition. However, the generalized Bohr frequency condition used there does not have a clearly presented classical counterpart.
@@
One is about a symmetry in the equation. Eq. (136) is invariant under the transformation of Hamiltonians $(H_{a_{k}})_{l_{1}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}}\to(H^{\prime}_{a_{k}})_{l_{1}~\!\!l_{2}~\!\!\cdots~\!\!l_{n}}$ whose values given by eq. (129) are transformed as
@@
$~~~~\times\prod_{r=1}^{k-1}(h_{a_{r}})_{l_{r}~\!\!l_{k}}\prod_{r^{\prime}=k}^{n-2}(h_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~\!\!l_{k}}$ (144)
@@
The other is about a solution of eq. (136), in the case that eq. (136) is regarded as a master equation. There is another solution such that
@@
$~~~~\times\prod_{r=1}^{k-1}(\tilde{h}_{a_{r}})_{l_{r}~\!\!l_{k}}\prod_{r^{\prime}=k}^{n-2}(\tilde{h}_{a_{r^{\prime}}})_{l_{r^{\prime}+1}~\!\!l_{k}}$ (148)
@@ 4.2 Discretization of Jacobian
where sgn(P) is +1+1 and 1-1 for even and odd permutations among the set of subscripts $(a_{1},a_{2},\cdots,a_{n})$, respectively, based on +1+1 for $(a_{1},a_{2},\cdots,a_{n})=(1,2,\cdots,n)$, and each $a_{k}$ $(k=1,2,\cdots,n)$ runs from $1$ to $n$. Hence, we conjecture that the generalized Heisenberg’s equation of motion (136) can be a counterpart of the Nambu equation and the multiple commutator defined by
@@
The properties of the fundamental identity and the derivation rule in the generalized matrix mechanics suggest that the quantum version of the generator of generalized canonical transformations in Nambu mechanics may be restricted to conserved quantities. Therefore, in Nambu mechanical systems where transformations generated by time-dependent quantities play an important role, it may be challenging to properly define their quantum counterparts.
@@ 5 Conclusions and discussions
Inspired by the fact that Heisenberg’s matrix mechanics was derived from Hamiltonian mechanics under the guidance of the correspondence principle, we have investigated a mechanics involving discrete variables, starting from Nambu mechanics. As a result, we reconstructed a framework of mechanics for dynamical variables expressed in terms of generalized matrices, which can be regarded as an extension of matrix mechanics. The fundamental equation of this mechanics corresponds to a generalized version of Heisenberg’s equation of motion, formulated using an $n$-fold commutator. This equation can also be interpreted as the counterpart to the Nambu equation associated with canonical $n$-plets. Based on this interpretation, we reconfirmed that the multiple commutator involving generalized matrices can serve as a quantum or discrete version of the Nambu bracket, namely the Jacobian.
@@
In our generalized dynamics, the objective is to determine the values of Hamiltonians, i.e., $(h_{a_{k}})_{l_{u}~\!\!l_{i}}$ in eq. (129), taking the equation (136) as the master equation. The relationship between symmetries and conservation laws holds in a manner analogous to conventional matrix mechanics. That is, generalized matrices $C^{a}$ that commute with the Hamiltonians, i.e., $C^{a},H_{1},\cdots,H_{n-1}]=0$, are identified as conserved quantities, as understood from the equation (136). Furthermore, arbitrary constant normal matrices become conserved quantities because the relation $N_{1},\cdots,N_{n}]=0$ holds for any $n$-th power normal matrices $N_{b}$ ($b=1,\cdots,n$).
In quantum mechanics, the matrix elements in matrix mechanics are interpreted as probability amplitudes representing transitions between quantum states. At present, it is unclear whether the generalized matrix elements with $n$ indices admit a similar interpretation, or whether the extended version of matrix mechanics can be applied to real physical systems, although the generalized cubic matrix mechanics can describe physical systems in the conventional matrix mechanics, as shown in the Appendix B. This suggests the need for further investigation into the understanding and application of generalized matrices, beyond the framework of a mere extension of Heisenberg’s matrix mechanics or a counterpart to Nambu mechanics.
@@
Finally, generalized matrices may prove useful in formulating the quantum theory of extended objects. The action integral in string theory includes the area of the worldsheet, and this area is expressed using a two-dimensional Jacobian [26, 27]. Similarly, $p$-dimensional extended objects are described using $(p+1)$-dimensional Jacobians. By discretizing the Jacobian, it may become possible to microscopically formulate various branes including membrane [28], which could, in turn, be applicable to the formulation of M-theory [^9] Using Lie 3-algebra [29], an effective theory for multiple M2 branes called the BLG model has been constructed [30, 31, 32, 33]. See Ref. [34] for extensive reviews of $n$-nary algebra and its applications.. Recently, there have been attempts relating to many-index objects, for instance, to regularize the Nambu brackets using $2n$-index objects [35] and to construct gauge theories using a new type of non-associative algebra that includes both ordinary matrices and cubic matrices [36]. These developments suggest that there may still be room for various extensions and applications.